OpenStudy (anonymous):

For the function f(x)= 4x-2, what is the range of f(x) for the domain {-2,0,3}?

OpenStudy (bibby):

What are the respective y values for every x value?

OpenStudy (anonymous):

There is none, or at least I don't understand what you are asking me @bibby

OpenStudy (bibby):

this function relates an x and y together in the form of an ordered pair. for example if the function was y=2x, and the domain was {-5,1,6} \(y=2*-5=-10\) \(y=2*1=2\) \(y=2*6=12\) the range would be {-10,2,12}

OpenStudy (anonymous):

So in My question it would be: 

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

exactly

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

So then the range would be {-10,-2,10} @bibby

OpenStudy (bibby):

looks right

OpenStudy (anonymous):

Is there an order you need to put it in? Or is it just like that?

OpenStudy (bibby):

least->greatest As far as I know

OpenStudy (anonymous):

So that is the right answer (the one I just posted)

OpenStudy (bibby):

"looks right" -bibby, circa 2014

OpenStudy (anonymous):

Thank You @bibby !!!! :D :D

OpenStudy (bibby):

np np np :D

OpenStudy (anonymous):

Thanks for the medal :)